c= circumfernce d= diameter r= radius pie= 3.14 or 22/7
c= 44 meters
what is the radius
formula c=2*3.14*r
i do not know the answer

Answer by checkley71(8403)

44=2*(22/7)*R OR 44=(44/7)*R OR R=44(44/7) OR R=44*(7/44) OR R=7